Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O4b)
berni - 15:01 Thursday 11 April 2024 (63931) Print this report
Operator Report - Morning shift

The locking troubles of the previous shifts were still in present at the beginning of this shift.

The commissioning team worked all the morning to find the source of the problem; right at the end of the shift we could lock in LN3_SQZ (see next dedicated entry for details) and set Science mode.

 

ISYS

At 7:48 UTC, taking advantage of an unlock, the INJ team recovered the BPC position.

Images attached to this report
Comments to this report:
mwas - 18:14 Thursday 11 April 2024 (63934) Print this report

The ITF mode was not put this morning in Troubleshooting mode as it was last night. It would have been better to put it back to Troubleshooting when the activity restarted in the morning.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×